(5m-n+1) - (10-m-n)
5m-m= 4m?
5+10= 15?

Answer by rothauserc(4718)

(5m-n+1) - (10-m-n) = 5m-n+1 -10 +m +n
5m-n+1 -10 +m +n = 6m - 9 = 3*(2m-3)